Cos3A+Cos3B+Cos3C+Cos3π=0, then the least value of sum of two angles is
 

Grade:12th pass

1 Answers

Arun
5 years ago

Dear student,

 

COS3B+COS3C+COS3A – 1 = 0

COS3B+COS3C+COS3A=1
COS3B+COS3C+COS3A-1=0
COS3(B+C)/2.COS3(B-C)/2-SIN^3A/2=0
SOLVING WE GET
COS3C/2.COS3B/2.COS3A/2=0
EITHER3C/2=pi
C= 2 pi/3     
 
Hence A+B = pi/3
 
Hence least value = pi/3